# AsynEIO: Asynchronous Monocular Event-Inertial Odometry Using Gaussian Process Regression

### v1

@@ III-C WNOJ Prior Residual
in which $\Delta t_{k}=t_{k+1}-k_{k}$, $\bm{\xi}_{k,k+1}=\log(\bm{T}_{k}^{-1}\bm{T}_{k+1})^{\vee}$, $\bm{\mathcal{J}}_{k,k+1}=\bm{\mathcal{J}}(\bm{\xi}_{k,k+1})$, and $\bm{\varpi}_{k}=\bm{\varpi}(t_{k})$. Notice that $\log(\bullet)$ and $(\bullet)^{\vee}$ are the inverse maps of $\exp(\bullet)$ and $(\bullet)^{\land}$, respectively. We can compute the covariance matrix for prior residual from (15). Furthermore, the detailed Jacobians of prior residuals are listed in Appendix-C.
IV GP-Inertial Factor
The GPIF is directly built upon the continuous-time trajectory with the GP-based motion prior. The relationships between GP-based motion states and raw inertial measurements are derived as follows. Subsequently, the residuals and Jacobians are provided analytically.
IV-A Inertial Measurement Model
@@
where $Q_{a},Q_{g},Q_{bg},Q_{ba}$ are covariance matrices.
IV-B GP-Inertial Residual and Jacobian
Before describing the GPIF, we need first know the kinematic state at an arbitrary timestamp. To query the trajectory at timestamp $\tau$ between $t_{k}$ and $t_{k+1}$, we firstly interpolate the local variable $\bm{\gamma}_{k}(\tau)$ with
@@ IV-B GP-Inertial Residual and Jacobian
where the derivations of $\frac{\partial\bm{x}(\tau)}{\partial\bm{x}(t_{k})}$ can be found in the Appendix-D. Here, the nonzero components of $\frac{\partial\bm{e}(\tau)}{\partial\bm{x}(\tau)}$ are listed as follows:
@@
V GP-Preintegration
The GPP provides a different insight to utilize GP regression framework to tightly fuse inertial and other asynchronous measurements. Rather than conducting the kinematic prior, the GPP assumes local inertial measurements as a group of independent GPs and instantiates them on-the-fly in a data-driven manner.
V-A GP Assumptions on Local Measurements
@@
in which $\bm{C}(t)$ is the rotation matrix of the body frame with respect to the inertial frame at the timestamp $t$, and $\bm{a}(t)$ is so called the proper acceleration of the body frame. From the definition of (17), we know that $\bm{a}(t)=\bm{C}_{bi}(\ddot{\bm{r}}_{i}^{bi}-\bm{g}_{i})$. As shown in (25) and (26), the local accelerations and angular velocities are expressed in a local body frame at the timestamp $t_{k}$. On this basis, these variables can be integrated linearly.
V-B GP-Preintegration on Induced Latent States
@@
[…]
@@
V-C Residual Terms for GP-Preintegration
@@ VI Event-driven VIO
Given the GPIF and GPP , we design an event-inertial odometry that can estimate motion trajectory from pure event streams and inertial measurements. Both GPIF and GPP, in conjunction with the discrete preintegration [46], are built into the same EIO system (as shown in Fig. 2). This designment allows us to switch to different inertial fusion methods and compare them fairly.
VI-A Event-Driven Frontend
To realize high-temporal resolution, we design an asynchronous event-driven frontend where sparse features are detected and tracked in an event-by-event manner, as depicted in Fig. 3. Although separate event-driven feature detecting and tracking methods have been proposed in previous literature, a whole event-driven frontend needs further design on strategies of screening and maintenance. Concretely, we leverage a registration table to record the unique identifier of each active event feature. The pixel location of each event feature can be infer from its index in the registration table. Then, the unique identifier is mapped to the corresponding feature trajectory and its independent tracker. When a event measurement is obtained, we first search in the registration table to determine if its neighbor region already exists activate event features. If there is at least one active event feature, we utilize the identifier-tracker map to find the corresponding tracker and feed the event measurement to it. The movement of active feature will be reported by the tracker and collected into the feature trajectory. Otherwise, we update the SAE and determine if a FA-Harris [47] corner feature can be detected at the pixel location of this event measurement.
Figure 3: Event-driven feature trajectory generation and management.
@@ VI-A Event-Driven Frontend
Result: Fts trajectory set $\bm{F}=\{\bm{F}_{i}|i\in\mathcal{M}\}$.
@@ 7             $\mathcal{P}_{k}\leftarrow\mathcal{M}(k)$;
8             Update tracker $\mathcal{T}_{k}$ with $\bm{e}_{j}$;
@@
15                   Set $-1$ in $\mathcal{G}$ at $\mathcal{T}_{k}.location$;
@@ 21                   $i\leftarrow i+1$;
24                   Create a new tracker $\mathcal{T}_{i}$, Update with $\bm{e}_{j}$;
@@ 26                   Create $\mathcal{P}_{i}$ with $\bm{F}_{i}$ and $\mathcal{T}_{i}$;
Algorithm 1 Event-Driven Visual Frontend
Since the registration table has the same width and height as the resolution of event cameras, it is high-efficient to access and modify elements in the registration table. The search of existing features only pays attention to a small patch within the registration table, and thus has a low time complexity. The small patch also reject the feature clustering and realize sparse features. Similarly, the detection only occurs at the location of the current event measurement and is thus efficient.
As the tracking of active features is triggered by subsequent events, a special case is that an active feature have not be triggered for a long time. For example, a corner feature is detected in noisy events by mistake, and thus it should be discarded promptly. To realize it, we assume the active feature should be triggered frequently within a maximum time threshold. We periodically delete the active features that have not been updated for more than the maximum time threshold. In addition, we also set a minimum time threshold to avoid the feature trajectory collecting a mass of redundant measurements. With patch search and time threshold mechanisms, the result frontend can obtain sparse and asynchronous feature tracking.
VI-B Asynchronous Repojection
@@
When a feature trajectory has collected enough measurements for triangulation, the induced projection will be utilized to triangulate and embed these measurements into the underlying factor graph. Given the triangulated landmark $\bm{l}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ and the event feature measurement $\bm{m}_{j}^{i}$ as defined in (32), the projection model can be given by
@@
where $\bm{K}$ indicates the intrinsic matrix, $d_{i}$ is the depth of this landmark in current camera frame, $\bm{T}_{bc}$ is the extrinsic parameter of the event camera w.r.t the IMU, which can be solved by calibration. The kinematic state $\bm{T}(t_{i,j})$ can be queried using (7) and (IV-B).
Different from the interpolated projection factor, the induced projection factor has reshaped the continuous-time trajectory using current accessible inertial measurements. Intuitively, it is more credible to build visual projection with this inertial induced factor. This concept is similar to the induce GP in the latent state preintegration method, while we leverage the proposed direct inertial factor to accomplish it. In contrast, the discrete preintegration scheme produces a unique pseudo-observation until the next sampling state. As a result, the asynchronous visual projection factors between inter-sampling states can only be interpolated under the motion prior assumption, which may sometimes lead to overly approximate results. Generally, both the direct factor and latent state preintegration can enhance the asynchronous projection of our event-driven VIO.
VI-C Inverse Depth Parametrization
Figure 5: Inverse depth parametrization.
The inverse depth parametrization [48] is adopted to enhance the numerical stability and realize undelayed initialization. Different from the filter-based method in [48], we leverage this parametrization to optimize the GP-based factor graph and derive the analytical Jacobians. Furthermore, we define the directional vector in local frames rather than global frames, which is compatible with the optimization method. The (33) can be simplified as
@@ VI-C Inverse Depth Parametrization
We use the inverse depth $\rho$, the directional vector $\bm{\kappa}$ in the unit sphere of the first camera pose $\bm{T}_{1}$ observing this feature trajectory $\bm{F}_{i}$ to represent the landmark $\bm{l}_{i}$. As displayed in Fig. 5, the latter camera pose $\bm{T}_{2}$ should observe the ray $\bm{h}_{2}$ and have a reprojected pixel location $\bm{q}_{2}$. When the poses and landmark are estimated precisely, this reprojected pixel location should coincide with the visual tracking result $\tilde{\bm{q}}_{2}$ at the same camera pose. In fact, the direction of $\bm{h}_{2}$ can uniquely determine the $\bm{q}_{2}$. From the geometric constraints in Fig. 5, the projected directional vector can be given by
@@
where
$\bm{P}=\begin{bmatrix}\bm{1}^{3\times 3},\bm{0}^{3\times 1}\end{bmatrix}.$ (36)
Therefore, the projection function can be written as
@@
where
$(\bm{h}_{2}^{\dagger})^{\flat}=\begin{bmatrix}h_{x}/h_{z}\\ h_{y}/h_{z}\\ 1\end{bmatrix}.$ (38)
It means the infinity landmark or the pure rotation when the inverse depth $\rho$ is 00 for all time. In both situations, the pose and inverse depth can be optimized without ambiguity. Therefore, this parametrization permits immediate initialization of feature trajectories. The analytical Jacobians can be derived using the chain rule, as detailed in Appendix-E. Although inverse depth has a better numerical performance in optimization, it introduces more state variables (6-DoF first observed pose) in each projection factor. When the first observed pose occurs in inter-samples of the GP trajectory, it will increase further due to the GP interpolation (Two adjacent trajectory points with 36 dimensions). To alleviate this problem, we interpolate on the feature trajectory to find the closest sampling timestamp on the motion trajectory. The interpolated pixel position is then used to compute $\bm{\kappa}$. This process initializes each feature trajectory for inverse depth optimization and reduces 30 dimensions for each projection factor. Benefited from high-time resolution of visual frontend, the interpolation on feature trajectories normally has negligible deviations.
VI-D Overall Cost Function in Sliding-Window
Figure 6: Factor graph of two asynchronous fusing schemes. (a) is the factor graph of AsynEIO with sparse GP priors and GPIFs, and the asynchronous projection is realized by GP interpolation on the motion trajectory. (b) is factor graph with GPP factors, and the GPP allows to analytically preintegrate at arbitrary timestamps. Therefore, the GPP can also support the asynchronous projection. The dynamic marginalization and sliding-window optimization are adopted to reduce the computational consumption.
The primary advancement of GPIF is that it avoids the integral operations on raw measurements. For the micro-electro-mechanical systems (MEMS) IMU, these integral operations would introduce collective integral errors and deteriorate the estimation system after a fewer second. Instead, we establish the detailed relationship between the differential SE(3) states of the kinematic trajectory and the physical measurements obtained from the IMU. We find that the residual, Jacobians and covariance matrix for GPIF are elegant and suitable to be embedded into the GP-based factor graph. The overall cost for our event-driven VIO with GPIF is given by
@@ VI-D Overall Cost Function in Sliding-Window
$+\underbrace{\sum\|\bm{e}_{c}(t_{i,j})\|_{\bm{R}}^{2}}_{\bm{event\ cost}}+\underbrace{\sum\|\bm{e}_{g}(\tau)\|_{\bm{Q}_{g}}^{2}+\sum\|\bm{e}_{a}(\tau)\|_{\bm{Q}_{a}}^{2}}_{\bm{inertial\ cost}},$ (39)
where $\bm{R}$ is the measurement covariance matrix of the event camera. We can see that there is only measurement covariance required by (VI-D) if the direct inertial factor is applied. Compared with preintegration schemes, we reduce the computational consumption paid for the covariance propagation. Also, this propagation is unreliable when the inertial measurements are partly lost or even asynchronous.
@@
VI-E Event and Inertial Initialization

### v2

@@ III-C WNOJ Prior Residual
in which $\Delta t_{k}=t_{k+1}-k_{k}$, $\bm{\xi}_{k,k+1}=\log(\bm{T}_{k}^{-1}\bm{T}_{k+1})^{\vee}$, $\bm{\mathcal{J}}_{k,k+1}=\bm{\mathcal{J}}(\bm{\xi}_{k,k+1})$, and $\bm{\varpi}_{k}=\bm{\varpi}(t_{k})$. Notice that $\log(\bullet)$ and $(\bullet)^{\vee}$ are the inverse maps of $\exp(\bullet)$ and $(\bullet)^{\land}$, respectively. We can compute the covariance matrix for prior residual from (14). Furthermore, the detailed Jacobians of prior residuals are listed in Appendix-B.
III-D GP-Inertial Residual and Jacobian
@@
where $\bm{Q}_{a},\bm{Q}_{g},\bm{Q}_{bg},\bm{Q}_{ba}$ are covariance matrices.
Before applying the GPIF, we need first know the kinematic state at an arbitrary timestamp. To query the trajectory at timestamp $\tau$ between $t_{k}$ and $t_{k+1}$, we firstly interpolate the local variable $\bm{\gamma}_{k}(\tau)$ [28] with
@@
where the derivations of $\frac{\partial\bm{x}(\tau)}{\partial\bm{x}(t_{k})}$ can be found in the Appendix-C. Here, the nonzero components of $\frac{\partial\bm{e}(\tau)}{\partial\bm{x}(\tau)}$ are listed as follows:
@@
IV GP-Preintegration
The GPP provides a different insight to utilize GP regression framework to tightly fuse inertial and other asynchronous measurements. Rather than conducting the kinematic prior, the GPP assumes local inertial measurements as a group of independent GPs and instantiates them on-the-fly in a data-driven manner [39].
IV-A GP Assumptions on Local Measurements
@@
in which $\bm{C}(t)$ is the rotation matrix of the body frame with respect to the inertial frame at the timestamp $t$, and $\bm{a}(t)$ is the so-called proper acceleration of the body frame. From the definition of (16), we know that $\bm{a}(t)=\bm{C}_{bi}(\ddot{\bm{r}}_{i}^{bi}-\bm{g}_{i})$. As shown in (24) and (25), the local accelerations and angular velocities are expressed in a local body frame at the timestamp $t_{k}$. On this basis, these variables can be integrated linearly.
IV-B GP-Preintegration on Induced Latent States
@@
$\Delta\tilde{\bm{\phi}}_{k}(\tau)$ $=\int_{t_{k}}^{\tau}\tilde{\dot{\bm{\phi}}}_{k}(t)dt,$ (28)
$\Delta\tilde{\bm{\nu}}_{k}(\tau)$ $=\int_{t_{k}}^{\tau}\tilde{\bm{a}}_{k}(t)dt,$
$\Delta\tilde{\bm{r}}_{k}(\tau)$ $=\int_{t_{k}}^{\tau}\int_{t_{k}}^{t}\tilde{\bm{a}}_{k}(s)dsdt,$
@@
IV-C Residual Terms for GP-Preintegration
@@ V Event-driven VIO
Given the GPIF and GPP, we design an event-inertial odometry that can estimate motion trajectory from pure event streams and inertial measurements. Both GPIF and GPP, in conjunction with their variants, are built into the same EIO system (as shown in Fig. 2). This design allows us to switch to different inertial fusion methods and compare them fairly.
Figure 3: Event-driven feature trajectory generation and management. The registration table (left) marks the active pixels patches, which limits the detection and tracking regions and enables an event-driven front-end (right). All potential event features (labeled as Discarded Fts) within the active Tracker Patch will be abandoned.
V-A Event-Driven Front-end
@@
Result: Feature trajectory set $\bm{F}=\{\bm{F}_{i}|i\in\mathcal{M}\}$.
@@ 7    $\mathcal{P}_{k}\leftarrow\mathcal{M}(k)$;
8      Tracking: Update tracker $\mathcal{T}_{k}$ with $\bm{e}_{j}$;
@@
15         Set 1-1 in $\mathcal{G}$ at $\mathcal{T}_{k}.location$;
@@ 21       $i\leftarrow i+1$;
23         Create a tracker $\mathcal{T}_{i}$, Update with $\bm{e}_{j}$;
@@ 25         Create $\mathcal{P}_{i}$ with $\bm{F}_{i}$ and $\mathcal{T}_{i}$;
Algorithm 1 Event-Driven Visual Front-end
To realize high-temporal resolution, we design an asynchronous event-driven front-end where sparse features are detected and tracked in an event-by-event manner, as depicted in Fig. 3. Although separate event-driven feature detecting and tracking methods have been proposed in previous literature, a whole event-driven front-end needs further design on strategies of screening and maintenance.
In AsynEIO, The event-driven front-end is abstracted as several basic operators: 1) Search, 2) Query, 3) Registration, 4) Tracking. Each reported event first invokes the Search operator to determine whether there already exists an active tracker within its neighbor (labeled as orange solid boxes in Fig. 3). The Query operator try to detect a new feature within the inactive zone, while the Tracking operator feed the reported event to the active tracker to update the feature trajectory (as visualized in Fig. 4). New created trackers and feature trajectories are recorded using the Registration operator. All events located within the active Tracker Patch (labeled as red dashed boxes) will not be queried, even if they are exact corner features, such as regions in green circles of Fig. 3. A registration table is designed to ensure that all operators only access or modify within small patches. Each active tracker is triggered by only a small fraction of events (within the Active Tracker). The event coordinates strictly constrain the pixel area of Query operators, which is more efficient than conventional feature detection in intensity images.
@@
As shown in Alg. 1, the registration table $\mathcal{G}$ records the unique identifier $i$ of new event features. The pixel location of each event feature can be infer from its index in the registration table. Then, the unique identifier is mapped to the corresponding feature trajectory $\bm{F}_{i}$ and its independent tracker $\mathcal{T}_{i}$ (combined into a pair variable $\mathcal{P}_{i}$). When an event measurement $\bm{e}_{j}$ is obtained, we first update the Surface of Active Events (SAE) $\mathcal{S}$ and Search the registration table $\mathcal{G}$ to determine whether active event features already exist in its neighboring region (Line 4-5). If there is at least one active event feature, we utilize the identifier-tracker map $\mathcal{M}$ to find the corresponding tracker and feed the event measurement to it (Line 6-8). The movement of active feature will be reported by the tracker and collected into the feature trajectory (Line 11). The searched unique identifier $k$ in the registration table will also be moved to the corresponding location (Line 12). Otherwise, we Query the SAE $\mathcal{S}$ to determine if a FA-Harris [47] corner feature can be detected at the pixel location of $\bm{e}_{j}$ (Line 16-17). Once a new event feature is reported, an individual tracker $\mathcal{T}_{i}$ [22] is created and added in the set $\Phi$ (Line 18-23) for subsequent tracking. Accordingly, it is registered in the registration table $\mathcal{G}$ (Line 24).
Since the active tracker is triggered by asynchronous events, the inactive trackers should be removed promptly. To realize it, we assume the active tracker should be triggered frequently within a maximum time threshold. We periodically delete active features that have not been updated beyond the maximum time threshold (Line 13-15). In addition, we also set a minimum time threshold to avoid the feature trajectory collecting a mass of redundant measurements (Line 9-12). With the registration maintenance and time-threshold screening mechanisms, the result front-end can obtain sparse and asynchronous feature tracking.
V-B Asynchronous Repojection
Figure 5: Inverse depth parametrization.
@@ V-B Asynchronous Repojection
When a feature trajectory has collected enough measurements for triangulation, the induced trajectory will be utilized to triangulate and embed these measurements into the underlying factor graph. Given the triangulated landmark $\bm{l}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ and the event feature measurement $\bm{m}_{j}^{i}$ as defined in (31), the projection model can be given by
@@
where $\bm{K}$ indicates the intrinsic matrix, $d_{i}$ is the depth of this landmark in current camera frame, $\bm{T}_{bc}$ represents the extrinsic transformation matrix between the event camera and IMU. The kinematic state $\bm{T}(t_{i,j})$ can be queried using (7) and (III-D) for the GPIF scheme. In contrast, the GPP can analytically infer $\bm{T}(t_{i,j})$ from induced latent states using linear operators, as shown in (25) and (28).
The inverse depth parametrization [48] is adopted to enhance the numerical stability and realize undelayed initialization. Simplify (32) as
@@
Define the inverse depth $\rho$ and the directional vector $\bm{\kappa}$ in unit sphere of the first observed camera pose $\bm{T}_{1}$ to represent the landmark $\bm{l}_{i}$. As displayed in Fig. 5, the latter camera pose $\bm{T}_{2}$ should observe the ray $\bm{h}_{2}$ and have a reprojected pixel location $\bm{q}_{2}$. When the poses and landmark are estimated precisely, this reprojected pixel location should coincide with the visual tracking result $\tilde{\bm{q}}_{2}$ at the same camera pose. According to the geometric constraints in Fig. 5, the projected directional vector can be given by
@@
where $\bm{P}=\begin{bmatrix}\bm{1}^{3\times 3},\bm{0}^{3\times 1}\end{bmatrix}$. Therefore, the projection function can be written as
@@
where $(\bm{h}_{2}^{\dagger})^{\flat}=[h_{x}/h_{z},h_{y}/h_{z},1]^{\top}$. The infinity landmark or the pure rotation can be represented by zero inverse depth $\rho=0$ for all time. In both cases, the pose and inverse depth can be optimized unambiguously, allowing for the immediate initialization of feature trajectories. Although inverse depth offers immediate initialization, it introduces additional state variables (6-DoF for the first observed pose) for each projection factor. When the first observed pose occurs between samples of the GP trajectory, the complexity increases further due to GP interpolation, involving two adjacent trajectory points with a total of 36 dimensions. To mitigate this issue, we interpolate along the feature trajectory to identify the nearest sampling timestamp on the motion trajectory for computing $\bm{\kappa}$. This process initializes each feature trajectory for inverse depth optimization and reduces 30 dimensions for each projection factor. Benefited from high-time resolution of visual front-end, the interpolation on feature trajectories normally has negligible deviations.
V-C Overall Cost Function in Sliding-Window
Figure 6: Factor graph of two main asynchronous fusing schemes. (a) Factor graph for GPIF, where asynchronous projection is achieved through GP interpolation on the motion trajectory. (b) Factor graph for GPP, which enables analytical preintegration at arbitrary timestamps to support asynchronous projection. The dynamic marginalization and sliding-window optimization are adopted to reduce the computational consumption.
The key innovation of GPIF lies in its ability to handle asynchronous IMU measurements while eliminating the need for preintegration operations. For the Micro-Electro-Mechanical Systems (MEMS) IMU, the preintegration would introduce collective errors and deteriorate the estimation results after a fewer second. Instead, the GPIF establishes detailed relationships between the differential SE(3) states of kinematic trajectories and the physical measurements obtained from IMUs. The residuals, Jacobians, and covariance matrices for GPIF can be elegantly formulated and seamlessly integrated into GP-based factor graphs. The overall cost for our event-driven VIO with GPIF is given by
@@ V-C Overall Cost Function in Sliding-Window
$+\underbrace{\sum\|\bm{e}_{c}(t_{i,j})\|_{\bm{R}}^{2}}_{\bm{event\ cost}}+\underbrace{\sum\|\bm{e}_{g}(\tau)\|_{\bm{Q}_{g}}^{2}+\sum\|\bm{e}_{a}(\tau)\|_{\bm{Q}_{a}}^{2}}_{\bm{GPIF\ cost}},$ (36)
where $\bm{R},\bm{Q}$ are the covariance matrices, the WNOJ prior $\bm{e}_{k}$ is defined as (15), the bias prior is expressed in (III-D), the event cost is denoted in (32), and the GPIF cost is given by (III-D). The estimated states of (V-C) include kinematic states $\bm{x}$ and landmarks $\bm{l}$ in the sliding-window. Specifically, the kinematic states $\bm{x}$ for GPIF consist of GP trajectory points $\bm{x}(t_{k})$ and IMU biases $\bm{b}_{a}(t_{k}),\bm{b}_{g}(t_{k})$}, as defined in (6) and (16). In addition, the event cost and GPIF cost in (V-C) leverage the interpolation formulas (III-D) and (7)-(11) to conduct asynchronous residuals. Compared with preintegration schemes, the inertial cost of GPIF directly applies the measurement covariance (i.e., $\bm{Q}_{g}$ and $\bm{Q}_{a}$), which avoids the computational consumption paid for the covariance propagation. Moreover, this propagation becomes unreliable when dealing with partial loss or asynchronous inertial measurements. Similarly, the summary cost for GPP can be defined as
$\mathop{\min}_{\bm{x},\bm{l}}\underbrace{\sum\|\bm{e}_{bg}(t_{k})\|_{\bm{Q}_{bg}}^{2}+\sum\|\bm{e}_{bg}(t_{k})\|_{\bm{Q}_{bg}}^{2}}_{\bm{bias\ prior}}+\underbrace{\sum\|\bm{e}_{c}(t_{i,j})\|_{\bm{R}}^{2}}_{\bm{event\ cost}}$
$+\underbrace{\sum\|\bm{e}_{\phi}(t_{k})\|_{\bm{Q}_{\phi}}^{2}+\sum\|\bm{e}_{\nu}(t_{k})\|_{\bm{Q}_{\nu}}^{2}+\sum\|\bm{e}_{r}(t_{k})\|_{\bm{Q}_{r}}^{2}}_{\bm{GPP\ cost}},$ (37)
in which the bias prior and event cost terms are the same as GPIF, and the GPP cost is defined as (29). Different from GPIF, The GPP has lower-dimension kinematic states including $\bm{T}(t_{k}),\bm{\nu}(t_{k}),\bm{b}_{a}(t_{k}),\bm{b}_{g}(t_{k})$}, where the interpolated state $\bm{x}(\tau)$ for asynchronous event residuals can be inferred from induced latent states and linear operators, as done in (25) and (28).
@@
V-D Event and Inertial Initialization